Break-Even Sales

Currently, the unit selling price of a product is $260, the unit variable cost is $210, and the total fixed costs are $640,000. A proposal is being evaluated to increase the unit selling price to $290.

a. Compute the current break-even sales (units).
units

b. Compute the anticipated break-even sales (units), assuming that the unit selling price is increased and all costs remain constant.
units
Answer #1

Correct Answer:

Requirement 1:

A

sales price

$               260.00

B

Less: Variable costs

$               210.00

C =A-B

Contribution margin

$                 50.00

D

Fixed costs

$       640,000.00

E = D/C

Break-even point (Units)

12800 Units

Requirement 2:

A

sales price

$               290.00

B

Less: Variable costs

$               210.00

C =A-B

Contribution margin

$                 80.00

D

Fixed costs

$       640,000.00

E = D/C

Break-even point (Units)

8000 Units

End of answer.

Please give a thumbs-up, it will be highly appreciated.

Thanks.